OpenStudy (anonymous):

Solve: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (anonymous):

-4x - 40 - 6 = -3x + 6 -4x - 46 = -3x + 6 -4x - 46 + 46 = -3x + 6 + 46 -4x = -3x + 52 -4x+3x = -3x+3x + 52 x = 52

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

−4(x + 10) − 6 = −3(x − 2) -4x - 40 - 6 = -3x +6 -46 - 6 = 4x - 3x x = -52